# Rethinking Discrete Speech Representation Tokens for Accent Generation

### v1

@@ 3.3 Accent, Speaker, and Phonetic ABX
We choose ABX as a model-free method to estimate the accessibility of accent information, as well as phone and speaker information in DSRTs.
In accent ABX, instances $a$ and $x$ share the same accent, while instance $b$ differs in accent. Triplets $(a,b,x)$ are constructed to share identical lexical content rather than phonetic context, as accent distinctions arise from accent-dependent realizations of the same words or lexical units, which may involve different phone sequences. We further require that $a$, $b$, and $x$ come from different speakers to avoid trivial similarity between $a$ and $x$ due to shared speaker identity, e.g.: a triplet $(a,b,x)$ consists of: $x=$ “water” spoken by a Scottish speaker1, $a=$ “water” by a Scottish speaker2, and $b=$ “water” by a Southern English speaker3.
We also adopt speaker ABX and phone ABX as contrastive evaluations. For detailed triplet selection criteria, see Appendix A.
In practice, words vary in both accent-induced pronunciation variation and speaker-specific pronunciation patterns. While standard ABX aggregates scores over all valid triplets, we adopt a word selection scheme to improve the sensitivity of accent ABX. We first select the 100 most frequent words in the training data, compute ABX scores for each (accent $A$ and $B$, word) combination, and retain the 10% with the lowest scores as the most accent-discriminative. The final accent ABX score is obtained by averaging their ABX scores on the test set. The combination selection is performed using continuous features from GenAID Zhong et al. (2025b), a supervised accent identification model fine-tuned from XLSR Babu et al. (2021), to ensure effective word selection and a fair comparison among DSRTs. See Appendix B for details.
(b) ABX evaluation results for information accessibility.
Figure 2: Accent, speaker, and phonetic information in different DSRTs across layers and representations.
(See Figure 4 in Appendix D for additional results from cross-accent VC.)
4 Experiments
4.1 Obtaining DSRTs
We extract continuous speech representations from various layers of HuBERT222hubert-large-ll60k, available here, HuBERT-ft333hubert-large-ls960-ft, available here and Whisper444whisper-meidum.en, available here. All three speech representation models share the same Encoder architecture (24 Transformer layers) and are trained on English-only data. Here, we focus on investigating how layer choice and ASR pretraining affect the information encoded in speech representations, and leave the investigation of how multilingual pretraining or larger-scale pretraining data/model to future work.
The extracted representations are then discretised using RepCodec [^5] https://github.com/mct10/RepCodec, which we train on the train-clean-100 subset of LibriSpeech [^6] https://www.openslr.org/12 following similar practice in Huang et al. (2024) and Zhang et al. (2025), with various codebook sizes. This subset contains 100 hours of speech with "accents closer to US English" Panayotov et al. (2015). RepCodec is trained for 200,000 steps with the same hyperparameters in Huang et al. (2024).
4.2 Evaluation Dataset
To evaluate information recoverability and accessibility with respect to accent, we use the VCTK corpus [^7] https://datashare.ed.ac.uk/handle/10283/3443, which provides relatively broad coverage of native English accents. Based on the provided country labels and detailed region descriptions, we manually group speakers into 13 accent regions, as some coarse country-level labels (e.g. “US”, “England”) conflate multiple distinct accent varieties.
We withhold 3–4 speakers from each accent region (50 speakers in total, with balanced gender and accent distribution) for testing. For the remaining speakers, we use 4 accent regions that have sufficient speakers (45 speakers in total) for training unit-to-speech HiFiGAN in Section 3.2 and for selecting accent-discriminative words in Section 3.3. Out of the 13 accent regions, 4 are seen in HiFiGAN training and accent ABX word selection, 5 are likely seen by RepCodec in LibriSpeech, and the remaining 4 are unseen at any stage, allowing for analysis on the generalisability of DSRTs across accents. See Appendix C for details of accent grouping and speaker splits.
Figure 3: Accent, speaker, and phonetic information recoverability in different DSRTs across codebook sizes.
(See Figure 5 in Appendix D for additional results from cross-accent VC and ABX.)
4.3 Evaluating Information Recoverability
To evaluate information recoverability, we train unit-to-speech HiFiGAN models [^8] https://github.com/facebookresearch/speech-resynthesis for 100,000 steps with the same hyperparameters as Polyak et al. (2021), using 4 accent regions (45 speakers), as described in Section 4.2. During inference, we perform cross-accent VC using source (src) DSRTs from 12 accent regions (46 speakers, excluding Southern English), together with speaker IDs from 4 Southern English speakers as target (tgt) voices. We fix the target voices to one accent region to control for target-side accent variability, enabling a clearer analysis of how accent information is preserved in the src DSRTs.
For objective evaluation of the converted speech, we adopt metrics recommended by Zhong et al. (2025a). (1) Accent similarity: We calculate cosine similarity of accent embeddings (Accent COS-SIM) extracted from GenAID [^9] https://github.com/jzmzhong/GenAID Zhong et al. (2025b). GenAID is trained with explicit speaker-accent disentanglement, preserving little speaker information in the extracted accent embeddings. (2) Speaker similarity: We calculate cosine similarity of speaker embeddings (Speaker COS-SIM) extracted from WavLM101010wavlm-base-plus-sv, available here Chen et al. (2022). These speaker embeddings encode some accent information, evidenced by improved performance in AID when transfer learned from SV models Zuluaga-Gomez et al. (2023). (3) Phonetic similarity: We calculate Jensen-Shannon distance between aligned PPGs [^11] https://github.com/interactiveaudiolab/ppgs (PPG Distance), extracted from a phone recognition model Churchwell et al. (2024). It is worth noting that PPG distance, primarily designed for pronunciation distance, is sensitive to accent information, as accent can be characterised partially as different realisations of the same phoneme. (4) Intelligibility: We calculate Word Error Rate (WER) using whisper-medium.en transcriptions, compared against ground-truth transcriptions. Despite its evident accent bias Jahan et al. (2025), WER is included here due to its broad adoption in previous DSRT evaluation frameworks to reflect how DSRTs are commonly selected.
For Accent COS-SIM, Speaker COS-SIM, and PPG distance, we use the first 24 utterances per speaker, which is an elicitation paragraph shared across speakers. In this way, we allow for similarity/distance calculation to either source speech that provides DSRTs or target speech that provides the speaker ID of the same content, removing the influence of content on these metrics. For WER, we randomly choose 24 utterances per speaker to cover diverse contents.
For visualisation, we annotate each metric with a practical lower bound (chance-level performance) and a practical upper bound (best possible performance), which anchor the visual scale and facilitate interpretation of values. For Accent COS-SIM, PPG, and WER, the lower bound measures the source speaker ground truth (GT) with respect to target speaker GT, and the upper bound measures the source speaker copy-synthesis with respect to source speaker GT. For Speaker COS-SIM, the lower bound measures similarity between source speaker GT and target speaker GT, while the upper bound measures similarity between target speaker copy-synthesis and target speaker GT.
4.4 Evaluating Information Accessibility
We follow the same training–test partition as in HiFiGAN experiments for simplicity. Accent ABX is conducted using this partition, whereas phone and speaker ABX are computed only on the test split, as they do not require a training phase. Test experiments used all utterances in the split, except for the first 24 utterances to avoid identical sentences across all train and test speakers, which could bias the ABX scores toward lexical-context matching instead of reflecting true accent or speaker discrimination. Phone-level alignment (phone boundaries and labels) required by phone ABX, were obtained through the Montreal Forced Aligner with the English Arpabet dictionary McAuliffe and Sonderegger (2024); Gorman et al. (2011).
Token type
Representation
& Layer
Codebook
size
A-SIM
A-SIM
S-SIM
[…]
S-SIM
[…]
Vevo Content HuBERT L18 32 0.8143 0.9369 0.6817 0.9651 0.1718 0.0824
Table 1: Results of proposed content and content-accent tokens, compared with content and content-style tokens from Vevo (Zhang et al., 2025). Differences in DSRT design choices are marked in red, with best performances marked in bold. *: For content tokens, we want little accent information recovered from DSRTs, i.e. low A-SIM (src) and high A-SIM (tar). For content-accent or content-style tokens, we want the opposite.
5 Results
5.1 Layer Choice Matters for Accent
5.2 ASR Supervision Removes Accent Information
5.3 Reducing Codebook Size Provides Limited Disentanglement
We then vary codebook sizes for three selected layers: HuBERT L18 (a common choice, as in Vevo Zhang et al. (2025)), HuBERT-ft L18 (a layer with little accent information while retaining high phonetic information), and HuBERT L9 (a layer with most accent information), shown in Figure 3. (1) Focusing on accent recoverability, varying codebook sizes has a markedly smaller effect than changing representation models or layers. For each layer, increasing codebook sizes from 32 to 2048 consistently improves accent and phonetic recoverability across all inspected DSRTs. However, the recoverability gap induced by codebook sizes variation within a layer is substantially smaller than that induced by changing layers at a fixed codebook sizes. (2) Speaker and phonetic recoverability also increase with larger codebook sizes. Taking HuBERT DSRTs as an example, Accent COS-SIM drops sharply as codebook sizes decreases from 1024 to 32, while PPG distance also rises 20.4%. Despite speaker identity being partially suppressed by the injected speaker ID during resynthesis, speaker recoverability also declines. The parallel degradation across accent, speaker, and phonetic metrics suggests minimal disentanglement when reducing codebook size.
5.4 Limitations of Commonly Used DSRTs in Speech Generation
Based on the above findings, we highlight the limitations of commonly used DSRTs here. (1) Vevo’s choice of using DSTRs based on HuBERT L18 with codebook size 32 as content tokens and codebook size 8,192 as content-style tokens Zhang et al. (2025) is suboptimal. Significant accent information is likely already lost and unrecoverable in HuBERT L18 - leading to insufficient accent information encoded in content-style tokens. Since naive reduction of codebook size cannot remove accent without harming phonetic information, content tokens achieve little accent information at the cost of intelligibility in resynthesised speech. (2) ASR-supervised DSRTs, adopted by Du et al. (2024), exhibit lower peak accent and phonetic recoverability than HuBERT DSRTs. Similar to content-style tokens, ASR-supervised DSRTs cannot encode accent sufficiently.
5.5 Proposed Content and Content-Accent Tokens
For better accent generation and control, we propose content-accent and content tokens for accent-preserving and accent-adaptive VC, respectively. A comparison of proposed tokens and Vevo tokens is presented in Table 1. Our proposed content/content-accent tokens achieve better performance than Vevo’s content/content-style tokens across all metrics, enabling better control of accent in speech generation. Demos of the converted speech will be available at [Link redacted].
6 Discussion
We observed a systematic delay of the optimal accessible layer relative to the recoverability peak. As discussed in (Yeh and Tang, 2024), earlier layers retain more complete acoustic information that can be effectively extracted by a strong HiFiGAN vocoder, whereas later layers gain phonetic discriminability from ASR-related objectives,
Across both metrics, accent information follows a distinct yet intermediate trend between phonetic and speaker information. This aligns with linguistic intuition: accent is partially realised through systematic phonetic variation and partially through speaker-specific pronunciation traits.
Finally, our findings provide potential explanations for the hallucinated accents by ZS-TTS systems Zhong et al. (2025b). Many ZS-TTS systems rely on representations (often deep or supervised layers) where accent recoverability is already attenuated, encouraging the model to guess accent patterns in inference or default to mainstream accents in training data.
7 Conclusion
Limitations
Although the framework we proposed guided us to develop a pair of DSRTs achieving good control over accent in VC, we acknowledge several limitations. First, we did not evaluate the proposed DSRTs in ZS-TTS settings, as training ZS-TTS systems requires substantial computational resources. Future work will extend our evaluation to ZS-TTS to assess the generalisability of our framework beyond voice conversion. Second, the current Accent ABX test shows low accent accessibility across all evaluated DSRTs. A supervised accent representation (GenAID Zhong et al. (2025b)) achieves much lower ABX error rate (9.87%), indicating the validity of the metric. More powerful probing models are needed to better reveal subtle accent information in SSL representations and DSRTs without explicit accent-discriminative training, so as to provide more precise guidance for accent applications. Better design of DSRTs is also needed to make accent information more accessible, and thus easier to learn by downstream speech generation models. Third, our experiments are limited to English SSL models and the VCTK dataset, which contains only L1 English speaker accents. We plan to extend the framework to multilingual representations and datasets involving L2 English accents and non-English languages to improve regional and language coverage. Finally, we focus on objective metrics to evaluate DSRTs and their effectiveness in VC. We haven’t included subjective evaluation and listening test, as Zhong et al. (2025a) have shown the effectiveness of the adopted objective metrics on evaluating accent generation, correlating with subjective listening tests results.
Broader Impacts and Ethical Statements
This work improves the transparency and controllability of accent information in discrete speech representation tokens (DSRTs). The proposed framework may help mitigate accent hallucination and unintended accent drift, benefiting speakers with underrepresented accents and potentially supporting applications such as speech therapy through accent- and voice-preserving speech generation. Our results further suggest that accent control can be partially achieved via informed representation and layer selection, potentially reducing reliance on additional large-scale model training.
Although this work is intended for research use, enhanced accent controllability in speech generation could be misused for voice impersonation if deployed without safeguards. If extended to real-world or unseen speakers, practical systems should incorporate mechanisms to ensure speaker approval and to detect synthesized speech. We emphasise that our framework is designed to analyse and expose limitations of DSRTs rather than to enable unrestricted accent manipulation.
This study exclusively uses existing publicly available datasets and pretrained models for research purposes, without collecting new data or releasing derived datasets. All data and pre-trained models are used under the following licenses: Creative Commons BY 4.0 License, Creative Commons BY-NC-ND 4.0 License, Apache License 2.0, and MIT License.
Acknowledgments
This work was supported in part by the UKRI AI Centre for Doctoral Training in Responsible and Trustworthy in-the-world Natural Language Processing (Grant EP/Y030656/1), UKRI Centre for Doctoral Training in Natural Language Processing (Grant EP/S022481/1), School of Informatics, and School of Philosophy, Psychology & Language Sciences, the University of Edinburgh. We would like to thank Dr. Hao Tang and Dr. Tianzi Wang for useful discussion.
Appendix A ABX Conditions
@@
Phone phone phone, context speaker
Table 2: Different ABX conditions for selecting $(a,b,x)$ triplets to evaluate accent, speaker, and phonetic information accessibility.
: context refers to the previous and next phones.
[…]
Appendix B Selected ABX Accent-word Combination
After obtained a list of (accent $A$ and $B$, word) combinations with scores measured on GenAID features, we take the first p% of combination with lowest scores, which captures the most typical accent-specific pronunciation differences. A grid search on $p\%$ is performed over 2.5%, 5%, 10%, 15% and 20%, and we reported results at 10%, which is the lowest percentage where the score starts to show a converging trend.
We take the first 25 selected combination as an example, presented in Table 3.
Accent Word Accentb
@@ Appendix B Selected ABX Accent-word Combination
Table 3: An example of the 25 triplet of (Accent AA, Accent $B$, word) with lowest ABX error rates, selected based on GenAID Zhong et al. (2025b) features on VCTK, applying to evaluate DSRTs accent ABX error rate in the test.
The table reveals a few well-established phonetic dimensions that underlie accent discrimination, successfully captured by our data-driven method. First, rhoticity emerges as a dominant cue. Words such as first and work sharply distinguish Southern or Northern English from Scottish and Irish accents, reflecting the systematic presence of post-vocalic /r/ in Scottish and Irish English and its absence in Southern English. The table shows that our metric takes good advantage of using rhoticity difference.
Second, differences in vowel quality, particularly in stressed lexical vowels, play a central role. For instance, first and year involve the near-front vowel spaces, which are realized as long centralized vowels in Southern English but shift toward more open or fronted realizations in Scottish and Irish English.
Third, consonantal realization, especially of /t/, contributes substantially to discrimination. Words such as scottish and last contain environments where Scottish English strongly favors glottal or lenited realizations of /t/, in contrast to the more alveolar realizations found in Southern and Irish English.
Finally, patterns of vowel reduction and weak-form realization further differentiate accents. Function words such as however, other, and from show strong schwa reduction in Southern English, whereas Irish and Scottish English tend to preserve fuller vowel qualities. These differences in reduction strategies, particularly in unstressed positions, amplify accent contrasts and explain why relevant accent-word combiantion are picked in the table.
From the analysis above, we can see the accent-word combination selected to evaluate accent ABX has a strong grounding in phonetic feature. The low accent accessibility results shown in Figure 4 and 3 authentically reflect the lack of accessible accent encoding in current speech representations and DSRTs.
Appendix C Accent and Speaker Partition in Cross-Accent Voice Conversion
See Table 4.
Accent Group Accent Region Split Speaker IDs
Seen during HiFiGAN training & Accent ABX word selection Southern English Train/Valid p225, p226, p228, p229, p231, p232, p239, p240, p243, p250, p254, p257, p258
Test p268, p273, p274, p276
Northern English Train/Valid p227, p230, p233, p236, p244, p256, p259, p267, p269, p270, p278, p279
Test p277, p282, p286, p287
Scottish Train/Valid p234, p237, p241, p246, p247, p249, p252, p255, p260, p262, p263, p271, p272, p275, p281
Test p264, p265, p284, p285
Irish Train/Valid p245, p266, p283, p288, p295
Test p298, p313, p340, p364
Likely seen during RepCodec training Canadian Test p316, p317, p343, p363
American Northeast Test p315, p339, p360, p361
American Midwest Test p311, p333, p334, p341
American South Test p301, p308, p310, p345
American West Test p294, p299, p300, p318
Unseen accents Northern Irish Test p292, p293, p304, p351
South African Test p314, p323, p336, p347
Indian Test p248, p251, p376
Oceanian Test p326, p335, p374
Table 4: Accent grouping and train/test splits for VCTK speakers. Accents are grouped by whether they were seen during HiFiGAN training & Accent ABX word selection, likely seen during RepCodec training, or unseen during training.
Appendix D Full Evaluation Results
Due to page limits, we report additional results here. Accent COS-SIM (to tar), Speaker COS-SIM (to tar), and WER exhibit reverse/similar trend, compared to Accent COS-SIM (to src), Speaker COS-SIM (to src), and PPG Distance, and are thus omitted in Figure 2 and 3. Information accessibility evaluation results across codebook sizes are also reported here in Figure 3.

### v2

@@ 3.3 Accent, Speaker, and Phonetic ABX
We choose ABX as a model-free method to estimate the accessibility of accent information, as well as phone and speaker information in DSRTs. Table 1 lists the three different ABX conditions we use to evaluate accent, speaker, and phonetic information accessibility.
Table 1: Different ABX conditions for selecting $(a,b,x)$ triplets to evaluate accent, speaker, and phonetic information accessibility. : context refers to the previous and next phones.
@@
Phone phone context speaker
When constructing $(a,b,x)$ triplets, a condition can be described as ABX ON the category of interest, BY the category of controlled constant among triplets, ACROSS the category of free variables among triplets. The classic Phone ABX select triplets ON phone BY context, where 1) $x$ and $a$ share the same phone, with $b$ of a different phone, and 2) $x$, $a$, and $b$ share identical phonetic context, e.g. bit [bIt] vs beat [bi:t]. To evaluate how prominent phonetic information is, compared to other variables, such as speaker, most Phone ABX would also be conditioned ACROSS speaker, with $a$, $b$ coming from the same speaker, and $x$ from a different speaker, e.g. $x=$ [bIt] by speaker1, $a=$ [bIt] by speaker2, $b=$ [bi:t] by speaker2.
In the proposed Accent ABX, instances $a$ and $x$ share the same accent, while instance $b$ differs in accent. Triplets $(a,b,x)$ are constructed to share identical lexical content rather than phonetic context, as accent distinctions arise from accent-dependent realisations of the same words or lexical units, which may involve different phone sequences. We further require that $a$, $b$, and $x$ come from different speakers to avoid trivial similarity between $a$ and $x$ due to shared speaker identity, e.g. a triplet $(a,b,x)$ comprosing $x=$ ``water'' spoken by a Scottish speaker1, $a=$ ``water'' by a Scottish speaker2, and $b=$ ``water'' by a Southern English speaker3.
In practice, words uttered by different speakers differ due to both accent-induced pronunciation variation and speaker-specific pronunciation patterns. While standard ABX aggregates scores over all valid triplets, we adopt a word selection scheme to improve the sensitivity of Accent ABX. We first select the 100 most frequent words in the training data, compute ABX scores for each (accent $A$, accent $B$, word) combination, and retain the 10% with the lowest scores as the most accent-discriminative. The choice of most frequent words is to ensure there are enough triplets to sample from in Accent ABX, while the selection of most accent-discriminative combinations is to avoid conducting Accent ABX on words that are pronounced similarly in two distinct accents. The combination selection is performed using continuous features from GenAID [zhong2025accentbox], a supervised Accent Identification (AID) model finetuned from XLSR [babu2021xls], to ensure effective word selection and a fair comparison among DSRTs.
Table 2: Samples of most accent-discriminative (Accent AA, Accent $B$, Word) combinations selected, for Accent ABX.
Accent A Word Accent B
@@
Table 2 shows the first 25 selected combinations as an example. The table reveals a few well-established phonetic dimensions that underlie accent differences, successfully captured by our data-driven selection scheme. First, rhoticity emerges as a dominant cue. Words such as first and work sharply distinguish Southern or Northern English from Scottish and Irish accents, reflecting the systematic presence of post-vocalic [r] in Scottish and Irish English and its absence in Southern English. The table shows that our metric takes good advantage of using rhoticity difference. Second, differences in vowel quality, particularly in stressed lexical vowels, play a central role. For instance, first and year involve the near-front vowel spaces, which are realised as long centralised vowels in Southern English but shift toward more open or fronted realisations in Scottish and Irish English. Third, consonantal realisation, especially of [t], contributes substantially to discrimination. Words such as scottish and last contain environments where Scottish English strongly favours glottal or lenited realisations of [t], in contrast to the more alveolar realisations found in Southern and Irish English. Finally, patterns of vowel reduction and weak-form realisation further differentiate accents. Function words such as however, other, and from show strong schwa reduction in Southern English, whereas Irish and Scottish English tend to preserve fuller vowel qualities. These differences in reduction strategies, particularly in unstressed positions, amplify accent contrasts and explain why relevant accent-word combinations are picked in the table.
4 Experiments
4.1 Obtaining DSRTs
We extract continuous speech representations from various layers of HuBERT [^2] https://huggingface.co/facebook/hubert-large-ll60k, HuBERT-ft [^3] https://huggingface.co/facebook/hubert-large-ls960-ft and Whisper [^4] https://huggingface.co/openai/whisper-medium.en. All three speech representation models share the same Encoder architecture (24 Transformer layers) and are trained on English-only data. Here, we focus on investigating how layer choice and ASR pretraining affect the information encoded in speech representations, and leave the investigation of how multilingual pretraining or larger-scale pretraining data/model to future work.
The extracted representations are then discretised using RepCodec [^5] https://github.com/mct10/RepCodec, which we train on the train-clean-100 subset of LibriSpeech [^6] https://www.openslr.org/12, similar to [zhang2025vevo, huang2024repcodec], with various codebook sizes. This subset contains 100 hours of speech with "accents closer to US English" [panayotov2015librispeech]. RepCodec is trained for 200,000 steps with the same hyperparameters in [huang2024repcodec].
4.2 Evaluation Dataset
To evaluate information recoverability and accessibility with respect to accent, we use the VCTK corpus [^7] https://datashare.ed.ac.uk/handle/10283/3443, which provides relatively broad coverage of native English accents. Based on the provided country labels and detailed region descriptions, we manually group speakers into 13 accent regions, as some coarse country-level labels (e.g. ``US'', ``England'') conflate multiple distinct regional accent varieties.
We withhold 3–4 speakers from each accent region (50 speakers in total, with balanced gender and accent distribution) for testing. For the remaining speakers, we use 4 accent regions that have sufficient speakers (45 speakers in total) for training unit-to-speech HiFiGAN in Section 3.2 and for selecting accent-discriminative words in Section 3.3. Out of the 13 accent regions, 4 are seen in HiFiGAN training and Accent ABX word selection, 5 are likely seen by RepCodec in LibriSpeech given they are North American accents, and the remaining 4 are unseen at any stage, allowing for analysis on the generalisability of DSRTs across accents.
4.3 Evaluating Information Recoverability
To evaluate information recoverability, we train unit-to-speech HiFiGAN models [^8] https://github.com/facebookresearch/speech-resynthesis for 100,000 steps with the same hyperparameters as in [polyak2021speech], using 4 accent regions (45 speakers), as described in Section 4.2. During inference, we perform cross-accent VC using source (src) DSRTs from 12 accent regions (46 speakers, excluding Southern English), together with speaker IDs from 4 Southern English speakers as target (tgt) voices. We fix the target voices to one accent region to control for target-side accent variability, enabling a clearer analysis of how accent information is preserved in the source DSRTs.
For objective evaluation of the converted speech, we adopt metrics recommended by [zhong2025pairwise]. (1) Accent similarity: We calculate cosine similarity of accent embeddings (Accent COS-SIM or A-SIM) extracted from GenAID [^9] https://github.com/jzmzhong/GenAID [zhong2025accentbox]. GenAID is trained with explicit speaker-accent disentanglement, preserving little speaker information in the extracted accent embeddings. (2) Speaker similarity: We calculate cosine similarity of speaker embeddings (Speaker COS-SIM or S-SIM) extracted from WavLM [^10] https://huggingface.co/microsoft/wavlm-base-plus-sv [chen2022wavlm]. These speaker embeddings encode some accent information, evidenced by improved performance in AID when transfer learned from SV models [zuluagagomez2023commonaccent]. (3) Phonetic similarity: We calculate Jensen-Shannon distance between aligned PPGs [^11] https://github.com/interactiveaudiolab/ppgs (PPG Distance), extracted from a phone recognition model [churchwell2024ppgs]. It is worth noting that PPG distance, primarily designed for pronunciation distance, is sensitive to accent information, as accent can be characterised partially as different realisations of the same phoneme. (4) Intelligibility: We calculate Word Error Rate (WER) using whisper-medium.en transcriptions, compared against ground-truth transcriptions. Despite its evident accent bias [jahan2025unveiling], WER is included here due to its broad adoption in previous DSRT evaluation frameworks to reflect how DSRTs are commonly selected.
We use the first 24 utterances per speaker, which is an elicitation paragraph shared across speakers, to calculate Accent COS-SIM, Speaker COS-SIM, and PPG distance. In this way, we allow for similarity/distance calculation to either source speech that provides DSRTs or target speech that provides the speaker ID of the same content, removing the influence of content on these metrics. As for WER, we randomly choose 24 utterances per speaker to cover diverse content.
For clearer visualisation, we annotate each metric with a practical lower bound (chance-level performance) and a practical upper bound (best possible performance), which anchor the visual scale and facilitate interpretation of values. For Accent COS-SIM, PPG, and WER, the lower bound measures the source speaker ground truth with respect to target speaker ground truth, and the upper bound measures the source speaker copy-synthesis with respect to source speaker ground truth. For Speaker COS-SIM, the lower bound measures similarity between source speaker ground truth and target speaker ground truth, while the upper bound measures similarity between target speaker copy-synthesis and target speaker ground truth.
Finally, we select several DSRTs to evaluate their performance in both accent-preserving VC and accent-adaptive VC, respectively, and conduct subjective listening tests on the converted speech. Three sets of 5-point Similarity MOS (SMOS) listening tests are conducted, with 15 listeners for each listening test, recruited on Prolific [^12] https://www.prolific.com/ from the target accent regions. To evaluate accent similarity between the generated speech and the source speech (Accent SMOS (src)) in accent-preserving VC, we randomly sample 20 utterances per selected DSRT, with speaker pairs drawn randomly from all 44 possible pairs (mapping from 11 American Midwest, Scottish, and Oceanian speakers to 4 Southern English speakers, covering both seen and unseen source accents). This results in 300 ratings per system per source accent region. To evaluate accent similarity between the generated speech and the target accent reference clip (Accent SMOS (tgt)) in accent-adaptive VC, we randomly sample 30 utterances per selected DSRT, again using random speaker pairs from the same 44 possible mappings. This results in 450 ratings per system. To evaluate speaker similarity between the generated speech and the target speaker reference clip (Speaker SMOS (tgt)) in both accent-preserving and accent-adaptive VC, we randomly sample 15 utterances per selected DSRT from the same pool of 44 speaker pairs, yielding 225 ratings per system.
4.4 Evaluating Information Accessibility
We follow the same training–test partition as in HiFiGAN experiments for simplicity. Accent ABX is conducted using this partition, whereas phone and speaker ABX are computed only on the test split, as they do not require a training phase. Test experiments used all utterances in the split, except for the first 24 utterances to avoid identical sentences across all train and test speakers, which could bias the ABX scores toward lexical-context matching instead of reflecting true accent or speaker discrimination. Phone-level alignment (phone boundaries and labels) required by phone ABX, were obtained through the Montreal Forced Aligner with the English ARPABET dictionary [mfa_english_us_arpa_acoustic_2024, gorman2011prosodylab].
(b) ABX evaluation results for information accessibility.